# StreamDiT: Real-Time Streaming Text-to-Video Generation

### v3

@@ 4.1 Model Architecture
Time-Varying DiT: As shown in Sec. 3.1, our StreamDiT changes the scalar condition $t$ to a sequence $\tau$ in the model. This requires that the time condition should be separable in the frame dimension. We follow the standard adaLN DiT [Peebles2022DiT] architecture, where time embedding is used for scale and shift modulations, and modify it with varying time embedding. Specifically, we reshape the latent tensor to 3D: $[F,H,W]$, and apply time embedding along the first dimension. As shown in Fig. 3, we modify the adaLN DiT with varying time embedding. Input tokens are also corrupted with different levels of noise. This architecture matches the design of our StreamDiT.
@@
Window Attention: To make the DiT model more efficient, we adopt window attention [swintransformer] for all self-attention layers. Specifically, we partition a 3D latent with shape $[F,H,W]$ to non-overlapped windows with size $F_{w},H_{w},W_{w}]$, and apply masking when computing self-attention, such that a token can only see tokens within the same window. To ensure consistency across windows, we shift by half of the window size every other layer. As illustrated in Fig. 4, voxels at axis ends will be warped over to the beginnings in the shifting. Accumulatively, global token communication can broadcast to all tokens, while each computation is efficient with local attention. The complexity of window attention is only $\frac{F_{w}\times H_{w}\times W_{w}}{F\times H\times W}$ of full attention.
Other Components: Since we are targeting real-time applications, we chose a moderate size (4B) for our DiT. We reuse a temporal auto-encoder (TAE) in Movie Gen [moviegen] with compression rate 4 in temporal domain and 8 in spatial domain. The latent channel size is 8, with the consideration that a moderate latent space is easier to learn for small generation models. The trade-off is that with a smaller temporal compression rate, the model needs to generate more latent frames to reach the same FPS.
For text encoders, we adopt the same ones from Movie Gen [moviegen], including UL2 [tay2023ul2unifyinglanguagelearning], ByT5 [xue2022byt5tokenfreefuturepretrained], and MetaCLIP [xu2024demystifyingclipdata]. The text encoders run fast on GPU. We only run text encoders when the prompt is changed for streaming video generation. Therefore, its computation time is negligible.
@@ 4.2 Multistep Distillation
Sampling distillation is a key component to build real-time streaming video generation. Due to the modifications of frame partitioning and micro denoising steps made by StreamDiT, standard sampling distillation methods [salimans2021progressive, guideddistillation, heek2024multistepconsistencymodels] cannot be applied. Thus, a customized sampling distillation for StreamDiT is needed.
@@ 5.1 Implementation Details
We finetune a T2V model [moviegen] with 4B parameters using the architecture introduced in Sec. 4.1. The latent size is $[16,64,64]$. With a $4\times,8\times,8\times]$ TAE, the base model generates 64 frames at 512p resolution.
Then we adapt the model for streaming video generation. Our StreamDiT training has three stages: task learning, task generalization, and quality fine-tuning. In the first stage, we use a small amount of high-quality video data (3K videos) with a large learning rate (1e-4) to adapt the original T2V into a video streaming model. The second stage involves further training on the pretraining dataset (2.6M videos) with a small learning rate (1e-5) to improve generalization for video streaming. In the final stage, we finetune the model on the high-quality video dataset with a small learning rate (1e-5) to optimize output quality. Each stage is trained with 10K iterations on 128 NVIDIA H100 GPUs.
To achieve real-time inference, we additionally apply the multistep distillation described in Sec. 4.2. We found a partitioning scheme $c=2,s=16,N=8$ is good for our multistep distillation considering both quality and efficiency, so we used it for teacher model inference. The distillation is also conducted on the 3K high-quality video dataset and 64 NVIDIA H100 GPUs for 10K iterations. After distillation, the micro step is reduced to 1, and the total number of sampling steps is 8.
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Quality
Score
@@ 5.2 Evaluation
We compare our method with ReuseDiffuse [gu2023reusediffuse] and FIFO-Diffusion [Kim2024FIFO] for streaming generation of long videos. To make a fair comparison, we implemented their methods on our base model, so they share the same visual quality. We select 50 prompts from the evaluation dataset of Movie Gen [moviegen] that are suitable for long videos. We adopt VBench [huang2023vbench] quality metrics for quantitative evaluation.
@@
In addition to VBench metrics, we also conduct human evaluations following the guidance in [moviegen]. A human evaluation compares two model results on the same prompts side-by-side. Annotators were asked to evaluate pairs of video samples to determine which one was superior or if they were of comparable quality along several axes. The evaluation criteria encompassed four aspects: overall quality, frame consistency, motion completeness, and motion naturalness. For these evaluations, we employ the same set of 50 prompts used in the VBench evaluation, generating 8-second videos at a resolution of 512p for each prompt. As shown in Fig. 5, our proposed method surpasses existing approaches across all evaluation metrics.
@@
Table 3: Ablation of mixed training. Chunk size 1 represents the Progressive AR Diffusion [xie2024progressiveautoregressivevideodiffusion], and chunk size 16 represents the original T2V without streaming. A mixed training with all chunk sizes show the best quality score.
@@ 5.3 Ablation Study
We conduct an ablation study to analyze the mixed training discussed in Sec. 3.2. Specifically, we train models with different mixing schemes as shown in Tab. 3. Chunk size 1 represents the Progressive AR Diffusion [xie2024progressiveautoregressivevideodiffusion] implemented on our model. Recall that our base model has 16 latent frames, so chunk size 16 is the basic T2V. After the models are trained, we generated videos using chunk size $c=1$, since this is the case covered by all models. Among them, we see that the mixture of all chunk sizes achieves the best quality score, although the inference is biased to chunk size 1 as being set in inference. Please note that we use a different inference scheme in Tab. 2, so the numbers of our model are different. The first one with chunk size 1 and no mixing achieves the second-best quality score, indicating an overfitting to this case.
@@ 5.4 Applications
Video-to-Video Streaming: Our model can also be applied to real-time video-to-video tasks. The video-to-video process is based on the noise addition and denoising strategy proposed in SDEdit [mengSDEditGuidedImage2022]. First, we add noise to the original input video and then denoise the noisy video with new prompt guidance. We achieve significant content editing while maintaining good temporal consistency. The visual results shown in Fig. 10 highlight the model’s powerful video editing capabilities. In this example, a pig in the input video is modified to a cat in the output video, while the background is preserved.
@@ 6 Conclusion and Limitations
Our StreamDiT model has only 4B parameters, which is moderate compared to some closed-source and open-source foundation models. The model capacity therefore limits the quality of basic T2V generation. As a result, we noticed artifacts in some of the generated videos, e.g. the Janus problem of the cat in our 5-minute video, and motion artifacts of the rally car. When applying our method on a larger model (30B Sec. 10 in Appendix) with a light training process, we found the quality improved a lot. From there, we can enhance the quality of small models via distillation for better real-time applications.
@@ 7 Acknowledgment
We would like to express our gratitude to Yen-Cheng Liu, Luxin Zhang, and Tao Xu for their valuable discussions and insightful contributions, which greatly enriched our work.
[…]
@@ 9 Design Choices and Training Details
Figure 15: High-quality videos generated by the 30B model, demonstrating the scalability of our method.
@@ 10 More Results
To evaluate the capabilities in generating higher-quality videos, we further fine-tuned the 30B model from Movie Gen [moviegen]. While the 30B model is not available for real-time generation at the present, it effectively supports the creation of exceptionally high-quality videos, as illustrated in Fig. 15. Our model consistently produces extended videos characterized by impressive visual quality, content coherence, and diverse scenes aligned accurately with provided text prompts.
@@ 11 Limitations
As described in Sec. 8, the effective context length of StreamDiT fundamentally depends on the base T2V model, and thus lacks long-term memory. When content falls outside the short-term memory window (i.e. context length) of StreamDiT, the associated information is likely to be lost. This can lead to issues such as identity mismatches in a person’s face, or background inconsistencies when the camera makes a full rotation. However, since StreamDiT is orthogonal to additional long-term memory mechanisms, it is possible to address this issue in the future by combining it with long-term memory architectures such as State Space Models like Mamba [gu2023mamba].

### v4

@@ 4.1 Model Architecture
Time-Varying DiT: As shown in Sec. 3.1, our StreamDiT changes the scalar condition $t$ to a sequence $\tau$ in the model. This requires that the time condition should be separable in the frame dimension. We follow the standard adaLN DiT [29] architecture, where time embedding is used for scale and shift modulations, and modify it with varying time embedding. Specifically, we reshape the latent tensor to 3D: $[F,H,W]$, and apply time embedding along the first dimension. As shown in Fig. 3, we modify the adaLN DiT with varying time embedding. Input tokens are also corrupted with different levels of noise. This architecture matches the design of our StreamDiT.
@@
Window Attention: To make the DiT model more efficient, we adopt window attention [22] for all self-attention layers. Specifically, we partition a 3D latent with shape $[F,H,W]$ to non-overlapped windows with size $F_{w},H_{w},W_{w}]$, and apply masking when computing self-attention, such that a token can only see tokens within the same window. To ensure consistency across windows, we shift by half of the window size every other layer. As illustrated in Fig. 4, voxels at axis ends will be warped over to the beginnings in the shifting. Accumulatively, global token communication can broadcast to all tokens, while each computation is efficient with local attention. The complexity of window attention is only $\frac{F_{w}\times H_{w}\times W_{w}}{F\times H\times W}$ of full attention.
Other Components: Since we are targeting real-time applications, we chose a moderate size (4B) for our DiT. We reuse a temporal auto-encoder (TAE) in Movie Gen [30] with compression rate 4 in temporal domain and 8 in spatial domain. The latent channel size is 8, with the consideration that a moderate latent space is easier to learn for small generation models. The trade-off is that with a smaller temporal compression rate, the model needs to generate more latent frames to reach the same FPS.
For text encoders, we adopt the same ones from Movie Gen [30], including UL2 [35], ByT5 [42], and MetaCLIP [40]. The text encoders run fast on GPU. We only run text encoders when the prompt is changed for streaming video generation. Therefore, its computation time is negligible.
@@ 4.2 Multistep Distillation
Sampling distillation is a key component to build real-time streaming video generation. Due to the modifications of frame partitioning and micro denoising steps made by StreamDiT, standard sampling distillation methods [32, 28, 12] cannot be applied. Thus, a customized sampling distillation for StreamDiT is needed.
@@ 5.1 Implementation Details
We finetune a T2V model [30] with 4B parameters using the architecture introduced in Sec. 4.1. The latent size is $[16,64,64]$. With a $4\times,8\times,8\times]$ TAE, the base model generates 64 frames at 512p resolution.
Then we adapt the model for streaming video generation. Our StreamDiT training has three stages: task learning, task generalization, and quality fine-tuning. In the first stage, we use a small amount of high-quality video data (3K videos) with a large learning rate (1e41e-4) to adapt the original T2V into a video streaming model. The second stage involves further training on the pretraining dataset (2.6M videos) with a small learning rate (1e51e-5) to improve generalization for video streaming. In the final stage, we finetune the model on the high-quality video dataset with a small learning rate (1e51e-5) to optimize output quality. Each stage is trained with 10K iterations on 128 NVIDIA H100 GPUs.
To achieve real-time inference, we additionally apply the multistep distillation described in Sec. 4.2. We found a partitioning scheme $c=2,s=16,N=8$ is good for our multistep distillation considering both quality and efficiency, so we used it for teacher model inference. The distillation is also conducted on the 3K high-quality video dataset and 64 NVIDIA H100 GPUs for 10K iterations. After distillation, the micro step is reduced to 1, and the total number of sampling steps is 8. For more training details, please refer to the Appendix.
Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Dynamic Degree Aesthetic Quality Quality Score
@@ 5.2 Evaluation
We compare our method with ReuseDiffuse [9] and FIFO-Diffusion [16] for streaming generation of long videos. To make a fair comparison, we implemented their methods on our base model, so they share the same visual quality. We select 50 prompts from the evaluation dataset of Movie Gen [30] that are suitable for long videos. We adopt VBench [15] quality metrics for quantitative evaluation.
@@
In addition to VBench metrics, we also conduct human evaluations following the guidance in [30]. A human evaluation compares two model results on the same prompts side-by-side. Annotators were asked to evaluate pairs of video samples to determine which one was superior or if they were of comparable quality along several axes. The evaluation criteria encompassed four aspects: overall quality, frame consistency, motion completeness, and motion naturalness. For these evaluations, we employ the same set of 50 prompts used in the VBench evaluation, generating 8-second videos at a resolution of 512p for each prompt. As shown in Fig. 5, our proposed method surpasses existing approaches across all evaluation metrics.
@@
Table 3: Ablation of mixed training. Chunk size 1 represents the Progressive AR Diffusion [39], and chunk size 16 represents the original T2V without streaming. A mixed training with all chunk sizes show the best quality score.
@@ 5.3 Ablation Study
We conduct an ablation study to analyze the mixed training discussed in Sec. 3.2. Specifically, we train models with different mixing schemes as shown in Tab. 3. Chunk size 1 represents the Progressive AR Diffusion [39] implemented on our model. Recall that our base model has 16 latent frames, so chunk size 16 is the basic T2V. After the models are trained, we generated videos using chunk size $c=1$, since this is the case covered by all models. Among them, we see that the mixture of all chunk sizes achieves the best quality score, although the inference is biased to chunk size 1 as being set in inference. Please note that we use a different inference scheme in Tab. 2, so the numbers of our model are different. The first one with chunk size 1 and no mixing achieves the second-best quality score, indicating an overfitting to this case.
@@ 5.4 Applications
Video-to-Video Streaming: Our model can also be applied to real-time video-to-video tasks. The video-to-video process is based on the noise addition and denoising strategy proposed in SDEdit [27]. First, we add noise to the original input video and then denoise the noisy video with new prompt guidance. We achieve significant content editing while maintaining good temporal consistency. The visual results shown in Fig. 10 highlight the model’s powerful video editing capabilities. In this example, a pig in the input video is modified to a cat in the output video, while the background is preserved.
@@ 6 Conclusion and Limitations
At the end, we discuss two limitations of our StreamDiT-4B, regarding model capacity and context length. StreamDiT-4B has a moderate model size, which limits the quality of basic T2V generation. As a result, we noticed artifacts in some of the generated videos.To verify the capacity limitation and the scalability of our method, we also applied StreamDiT to a 30B-parameter model with a light-weight training process. We found that StreamDiT-30B shows much better quality in terms of visual aesthetic, content, and low artifacts. Please refer to the Appendix and project website for the results. Another limitation of StreamDiT-4B is the short context length. Flaws may be observed whereby out-of-context objects exhibit altered visual appearance after they re-enter the video sequence. We anticipate that this limitation can be mitigated by employing an extended context window in conjunction with a key–value (KV) cache during inference, thereby enabling more efficient computational acceleration. Therefore, the two limitations are not fundamental to the StreamDiT method, and can be improved with future work.
@@ 7 Acknowledgment
We would like to express our sincere gratitude to Yen-Cheng Liu, Luxin Zhang, and Tao Xu for their valuable discussions and insightful contributions, which greatly enriched our work.
@@ 9 Design Choices and Training Details
Figure 15: High-quality long videos generated by StremDiT-30B, demonstrating the scalability of our method.
@@ 10 More Results
To evaluate the capabilities in generating higher-quality videos, we further fine-tuned the 30B model from Movie Gen [30]. While the 30B model is not available for real-time generation at the present, it effectively supports the creation of exceptionally high-quality videos, as illustrated in Fig. 15. Our model consistently produces extended videos characterized by impressive visual quality, content coherence, and diverse scenes aligned accurately with provided text prompts.
@@ 11 Limitations
As described in Sec. 8, the effective context length of StreamDiT fundamentally depends on the base T2V model, and thus lacks long-term memory. When content falls outside the short-term memory window (i.e. context length) of StreamDiT, the associated information is likely to be lost. This can lead to issues such as identity mismatches in a person’s face, or background inconsistencies when the camera makes a full rotation. However, since StreamDiT is orthogonal to additional long-term memory mechanisms, it is possible to address this issue in the future by combining it with long-term memory architectures such as State Space Models like Mamba [8].